Please help me with this algebra question. Thank you so much in advance!
I had sent this one in but forgot to put the question.
An equation and two ordered pairs have been given. Show each pair is a solution of the equation. Then if you can, graph the two pairs to determine another solution.
y = x - 2; (3,1),(-2,-4)

Answer by algebrahouse.com(1659)

y = x - 2; (3,1)
1 = 3 - 2 {substituted 3 in for x and 1 in for y}
1 = 1 {this is a true statement, therefore it is a solution}

y = x - 2; (-2,-4)
-4 = -2 - 2 {substituted -2 in for x and -4 in for y}
-4 = -4 {this is a true statement, therefore it is a solution}

Graphing:
- there are several ways to do this, so it is up to you
- an easy way to get another point without graphing is to
substitute any number in for x, other than 3 and -2 (those were used above)

For more help from me, visit: www.algebrahouse.com